# Simple homotopy types of even dimensional manifolds

### v2

@@ 10.3. Divisibility of class numbers of cyclotomic fields
The aim of this section will be to survey results on the divisibility of the class numbers $h_{m}$ and $h_{m}^{-}$. The most basic divisibility results are that, for $n\mid m$, we have $h_{n}\mid h_{m}$ [Wa97, p. 205] and $h_{n}^{-}\mid h_{m}^{-}$ [MM76, Lemma 5]. Motivated by Proposition 10.4, we will now pursue divisibility results of two distinct types. We will start by considering $\operatorname{odd}(h_{m}^{-})$, i.e. the unique odd integer $r$ such that $h_{m}^{-}=2^{k}r$ for some $k$, and we then consider the parity of $h_{m}^{-}$.
Recall the following theorem of Masley-Montgomery [MM76] (see also [Wa97, p. 205]). In anticipation of its application in the proofs of Theorems 11.1 and 11.2, we will separate out the case where $m$ is square-free.
@@
The proof of Proposition 10.5 is based on the fact that $h_{m}^{-}\to\infty$ as $m\to\infty$. We will now establish lower bounds on the growth rate of $h_{m}^{-}$, and hence on $h_{m}$ since $h_{m}\geq h_{m}^{-}$. Let $\varphi(m)$ denote Euler’s totient function.
@@
It is shown in [Wa97, Theorem 4.20] that $\log h_{m}^{-}/(\frac{1}{4}\varphi(m)\log m)\to 1$ as $m\to\infty$. This implies that $\log h_{m}^{-}\geq C_{0}\varphi(m)\log m$ for some $C_{0}>0$. The result now follows from the fact that $\varphi(m)\geq m/2\log\log m$ for $m$ sufficiently large [HW54, Theorem 328]. ∎
@@
We will prove this by tracing through Horie’s proof of [Ho89, Theorem 1]. Recall that an abelian field $K$ is a finite Galois extension $K/\mathbb{Q}$ with $\operatorname{Gal}(K/\mathbb{Q})$ abelian and we can assume that $K\subseteq\mathbb{C}$. For an abelian field $K$ with maximal real subfield $K^{+}$, let $h_{K}=|C(\mathcal{O}_{K})|$, $h_{K}^{+}=|C(\mathcal{O}_{K^{+}})|$ and $h_{K}^{-}=h_{K}/h_{K}^{+}$ (which is an integer). Let $\operatorname{disc}(K)$ denote the discriminant of a number field $K$.
@@
We will use that, if $L/K$ is an extension of abelian fields, then $\operatorname{odd}(h_{K}^{-})\mid\operatorname{odd}(h_{L}^{-})$ [Ho89, Lemma 1]. For each abelian field $K$, let $K^{\prime}$ denote the maximal subfield of $K$ with degree a power of 2. By the fundamental theorem of Galois theory and the fact that a finite abelian group $A$ has a subgroup of order $d$ for all $d\mid|A|$, we get that $|K^{\prime}/\mathbb{Q}|$ is the highest power of 2 dividing $|K/\mathbb{Q}|$. For an integer $n\geq 1$, let $\mathcal{A}_{n}=\{m\mid\operatorname{odd}(h_{m}^{-})\leq n\}$ and $\mathcal{B}_{n}=\{|\mathbb{Q}(\zeta_{m})^{\prime}/\mathbb{Q}|\mid n\in\mathcal{A}_{n}\}$.
We will begin by finding a bound for $\sup(\mathcal{B}_{n})$. Let $K=\mathbb{Q}(\zeta_{m})^{\prime}$ for some $m\in\mathcal{A}_{n}$. Then $\mathbb{Q}(\zeta_{m})/K$ is an extension of abelian fields and so $\operatorname{odd}(h_{K}^{-})\mid h_{m}^{-}$ and so $\operatorname{odd}(h_{K}^{-})\leq n$. Since $|K/\mathbb{Q}|$ is a power of 2, $h_{K}$ must be odd [Wa97] and so $h_{K}^{-}=\operatorname{odd}(h_{K}^{-})\leq n$. Furthermore, note that $K$ is imaginary unless $K=\mathbb{Q}$ [Ho89, p. 468].
@@
Let $K=\mathbb{Q}(\zeta_{m})^{\prime}$. Then $|\mathbb{Q}(\zeta_{m})/\mathbb{Q}|=2^{r}t$ for some $r\geq 0$ and $t$ odd, where $2^{r}=|K/\mathbb{Q}|\leq C\log n$ by the bound on $\log(\mathcal{B}_{n})$. Since $|\mathbb{Q}(\zeta_{m})/L|\mid|\mathbb{Q}(\zeta_{m})/\mathbb{Q}|$, we have that $2^{n(2)}\leq 2^{r}\leq C\log n$. Hence $h_{m}^{-}/\operatorname{odd}(h_{m}^{-})\leq h_{m}/\operatorname{odd}(h_{m})=2^{e_{N}^{(2)}}\leq(C\log n)^{C^{\prime}}$. Since $m\in\mathcal{A}_{n}$, we have $\operatorname{odd}(h_{m}^{-})\leq n$. This gives that $h_{m}^{-}\leq a(\log n)^{b}n$ for some constants $a,b>0$ and so, for any $\varepsilon>0$, we have that $h_{m}^{-}\leq an^{1+\varepsilon}$. Combining this with Proposition 10.6 gives that $\log(an^{1+\varepsilon})\geq C_{0}\frac{m\log m}{\log\log m}$ for some $C_{0}>0$, which implies that $\log n\geq C\frac{m\log m}{\log\log m}$ for some $C>0$.
@@
We will now consider the parity of $h_{m}$ and $h_{m}^{-}$. Whilst we will not explicitly make use of it, we will record the following basic observation which dates back to Kummer (see [Ha52, Satz 45], [Yo98, Remark 1]).
@@
Prior to the results of Ichimura–Nakajima, it was shown by Washington that $h_{p^{n}}/h_{p}$ is odd for $p=3,5$ [Wa75]. Note that both results of Washington and Ichimura–Nakajima all depend on Iwasawa theory. As far as we are aware, this is an essential ingredient in all known proofs that there exists an odd prime $p$ such that $h_{p^{n}}$ is odd for all $n\geq 1$.
@@ 10.4. Kernel groups of $\mathbb{Z}C_{m}$
The aim of this section will be to determine the involution on $D(\mathbb{Z}C_{m})$ which is induced by the involution on $C(\mathbb{Z}C_{m})$ (see Proposition 8.12).
We will begin with the following classical result due to Rim [Ri59, Theorem 6.24] (see also [CR87, Theorem 50.2]). Recall that, if $p$ is a prime and $\Gamma_{p}$ is the maximal order in $\mathbb{Q}C_{p}$ containing $\mathbb{Z}C_{p}$, then $\Gamma_{p}\cong\mathbb{Z}\times\mathbb{Z}[\zeta_{p}]$ and so $C(\Gamma_{p})\cong C(\mathbb{Z}[\zeta_{p}])$ since $\mathbb{Z}$ is a PID.
@@
We will now determine $D(\mathbb{Z}C_{m})$, as well as its involution, in the case where $m$ is square-free. This will be used in the proofs of Theorems 11.1 and 11.2. As usual, we will view an abelian group with involution as a $\mathbb{Z}C_{2}$-module.
@@
We will view this as a $\mathbb{Z}C_{2}$-module with the involution induced by the conjugation map
¯$\overline{\cdot}\colon(\mathbb{Z}[\zeta_{m}]/\pi_{m})^{\times}\to(\mathbb{Z}[\zeta_{m}]/\pi_{m})^{\times},\quad\zeta_{m}\mapsto\zeta_{m}^{-1}.$
Note that, if $p$ is prime, then $\mathbb{Z}[\zeta_{p}]/\pi_{p}\cong\mathbb{F}_{p}$. We can then see that $\Psi_{p}\colon\mathbb{Z}[\zeta_{p}]^{\times}\to\mathbb{F}_{p}^{\times}$ is surjective by considering the cyclotomic units $1+\zeta_{p}+\cdots+\zeta_{p}^{i-1}\in\mathbb{Z}[\zeta_{p}]^{\times}$ for $(i,p)=1$. In particular, $V_{p}=1$. We will now show how $D(\mathbb{Z}C_{m})$ is related to $V_{d}$ for $d\mid m$.
@@
First recall that $\Gamma=\bigoplus_{d\mid m}\mathbb{Z}[\zeta_{d}]$ is the unique maximal order in $\mathbb{Q}C_{m}$ which contains $\mathbb{Z}C_{m}$. This implies that there is a pullback square:
${\mathbb{Z}C_{m}}$${\Gamma}$${\mathbb{Z}C_{m}/m\Gamma}$${\Gamma/m\Gamma.}$$\scriptstyle{i_{2}}$$\scriptstyle{i_{1}}$$\scriptstyle{j_{2}}$$\scriptstyle{j_{1}}$
@@
is the $\mathbb{Z}C_{m}$-module with action $\lambda\cdot(x,y)=(i_{1}(\lambda)x,i_{2}(\lambda)y)$ for $\lambda\in\mathbb{Z}C_{m}$. We now claim that the conjugation map on $(\Gamma/m\Gamma)^{\times}$ induces the involution on $D(\mathbb{Z}C_{m})$. First note that, by Proposition 8.13, the involution on $D(\mathbb{Z}C_{m})$ is induced by the natural involution $x\mapsto x^{-1}$ on the idèle group $J(\mathbb{Q}C_{m})\subseteq\prod_{p}\mathbb{Q}_{p}C_{m}$. For all primes $p$, we have $\mathbb{Z}_{p}C_{m}\subseteq\Gamma_{p}\oplus(\mathbb{Z}C_{m}/m\Gamma)_{p}\subseteq\mathbb{Q}_{p}C_{m}$. If $M(u)=\mathbb{Z}C_{m}\alpha$ for $\alpha=(\alpha_{p})\in J(\mathbb{Q}C_{m})$, then [CR87, Exercise 53.1] implies that
@@
where $u_{p}\in\Gamma_{p}$ is any element such that $j_{2}(u_{p})=[u]\in(\Gamma/m\Gamma)_{p}$. By the same argument as in the proof of Proposition 10.1, the involution on $J(\mathbb{Q}C_{m})$ induces an involution on $J(\mathbb{Q}(\zeta_{d}))$ which coincides with the involution induced by conjugation. In particular, the involution maps $\alpha_{p}\mapsto(1,1)$ or $(\overline{u}_{p},1)$ where ¯$\overline{\,\cdot\,}\colon\Gamma_{p}\to\Gamma_{p}$ is induced by conjugation on $\Gamma$. In particular, this coincides with the involution induced by conjugation on $(\Gamma/m\Gamma)^{\times}$.
By [CR87, Lemma 50.7], $j_{1}$ can be replaced by the map $\alpha\colon(\mathbb{Z}C_{m}/m\mathbb{Z}C_{m})^{\times}\to(\Gamma/m\Gamma)^{\times}$. By [CR87, Lemma 50.8], $\operatorname{coker}(\alpha)\cong\bigoplus_{d\mid m}(\mathbb{Z}[\zeta_{d}]/\pi_{d})^{\times}$ and it follows from the proof that conjugation map on $(\Gamma/m\Gamma)^{\times}$ induces conjugation on $(\mathbb{Z}[\zeta_{d}]/\pi_{d})^{\times}$ for each $d\mid m$. If $\gamma\colon\Gamma^{\times}\to\operatorname{coker}(\alpha)$ is the map induced by $j_{2}$, then we obtain an exact sequence
[…]
@@
Since the involution on $D(\mathbb{Z}C_{m})$ is induced by conjugation on $(\Gamma/m\Gamma)^{\times}$, it follows that it restricts to $A_{i}$ where it acts via conjugation on the $\mathbb{Z}[\zeta_{d}]$. Hence, with respect to the involution, $A_{i}\leq D(\mathbb{Z}C_{m})$ is a $\mathbb{Z}C_{2}$-modules and the chain $A_{0}\leq\cdots\leq A_{n}$ is a chain of $\mathbb{Z}C_{2}$-modules. Under the abelian group isomorphism
@@
Let $m\geq 2$ be a square-free integer. We will now give a method for analysing the involution on $V_{m}=\operatorname{coker}(\Psi_{m})$. For a field $\mathbb{F}$ and $\alpha_{1},\dots,\alpha_{n}\in\mathbb{C}$, we will write $\mathbb{F}[\alpha_{1},\dots,\alpha_{n}]$ to denote $\mathbb{F}\otimes_{\mathbb{Z}}\mathbb{Z}[\alpha_{1},\dots,\alpha_{n}]$. Note that this may not be a field and so need not coincide with $\mathbb{F}(\alpha_{1},\dots,\alpha_{n})$.
First note that, as described above, we have that
@@
We will view this as a $\mathbb{Z}C_{2}$-module with the involution induced by the conjugation map
¯$\overline{\cdot}\colon\mathbb{F}_{p_{i}}[\zeta_{m/p_{i}}]^{\times}\to\mathbb{F}_{p_{i}}[\zeta_{m/p_{i}}]^{\times},\quad\zeta_{m/p_{i}}\mapsto\zeta_{m/p_{i}}^{-1}.$
@@
and $f(\overline{\alpha})=f(\alpha)^{-1}$. Since $f$ is surjective and induces the involution on $\widetilde{V}_{m}$, this implies that $\overline{x}=-x$ for all $x\in\widetilde{V}_{m}$, where we now write the inverse as $-x$ rather than $x^{-1}$ since $\widetilde{V}_{m}$ is an abelian group. ∎
We will now deduce the following, which is the main result of this section. Note that this is analogous to Proposition 10.4 which applied in the case of ideal class groups. Recall that, for an integer $m$, we let $\operatorname{odd}(m)$ denote the unique odd integer $r$ such that $m=2^{k}r$ for some $k$.
@@
In particular, $\prod_{d\mid m}\operatorname{odd}(|\widetilde{V}_{d}|)$ divides |$|\{x-\overline{x}\mid x\in D(\mathbb{Z}C_{m})\}|$.
@@
and $|2\cdot\widetilde{V}_{d_{n-i+1}}|$ divides |$|\{x-\overline{x}\mid x\in A_{i}\}/\{x-\overline{x}\mid x\in A_{i-1}\}|$ for all $1\leq i\leq n$. It now suffices to note that $\operatorname{odd}(|\widetilde{V}_{d_{n-i+1}}|)$ divides $|2\cdot\widetilde{V}_{d_{n-i+1}}|$ by the same argument as in the proof of Proposition 10.4. ∎
@@
To obtain a complete analogue of Proposition 10.4, it would be desirable to also obtain bounds on |$|\{x\in D(\mathbb{Z}C_{m})\mid x=-\overline{x}\}|$. However, unlike Proposition 10.4, the bounds we obtain in Proposition 10.17 are obtained by subquotients rather than just subgroups. We therefore cannot apply Lemma 9.3 (i) since the final map in the sequence need not be surjective in general (see Remark 9.4). As we shall see in Section 11, it is possible to circumvent the need for such bounds in the proof of Theorem 11.1.
We now conclude this section with a result which holds in the case that $m$ is not square-free. Firstly an analogue of Lemma 10.14 holds in the case that $m$ is a prime power, by Kervaire–Murthy [KM77, Theorem 1.2]. For brevity, we will not state this result here. We will instead make do with the following consequence of their result in the case that $p=2$ which will be used in the proof of Theorem 11.3. Note that $V_{2^{n+1}}$ is directly analogous to the $\mathbb{Z}C_{2}$-modules $V_{m}$ defined in Definition 10.13 when $m$ is square-free.
@@
[…]

### v3

@@ 10.3. Divisibility of class numbers of cyclotomic fields
The aim of this section is to survey results on the divisibility of the class numbers $h_{m}$ and $h_{m}^{-}$. The most basic divisibility results are that, for $n\mid m$, we have $h_{n}\mid h_{m}$ [Wa97, p. 205] and $h_{n}^{-}\mid h_{m}^{-}$ [MM76, Lemma 5]. Motivated by Proposition 10.5, we pursue divisibility results of two distinct types. We start by considering $\operatorname{odd}(h_{m}^{-})$, i.e. the unique odd integer $r$ such that $h_{m}^{-}=2^{k}r$ for some $k$, and we then consider the parity of $h_{m}^{-}$.
Recall the following theorem of Masley-Montgomery [MM76] (see also [Wa97, p. 205]). In anticipation of its application in the proof of Theorem 11.1, we separate out the case where $m$ is square-free.
@@
The proof of Proposition 10.6 is based on the fact that $h_{m}^{-}\to\infty$ as $m\to\infty$. We now establish lower bounds on the growth rate of $h_{m}^{-}$, and hence on $h_{m}$ since $h_{m}\geq h_{m}^{-}$. Let $\varphi(m)$ denote Euler’s totient function.
@@
It is shown in [Wa97, Theorem 4.20] that $\log h_{m}^{-}/(\frac{1}{4}\varphi(m)\log m)\to 1$ as $m\to\infty$. This implies that $\log h_{m}^{-}\geq C_{0}\varphi(m)\log m$ for some $C_{0}>0$. The result now follows from the fact that $\varphi(m)\geq m/(2\log\log m)$ for $m$ sufficiently large [HW54, Theorem 328]. ∎
@@
We prove this by tracing through Horie’s proof of [Ho89, Theorem 1]. An abelian field $K$ is a finite Galois extension $K/\mathbb{Q}$ with $\operatorname{Gal}(K/\mathbb{Q})$ abelian, and we can assume that $K\subseteq\mathbb{C}$. For an abelian field $K$ with maximal real subfield $K^{+}$, let $h_{K}=|C(\mathcal{O}_{K})|$, $h_{K}^{+}=|C(\mathcal{O}_{K^{+}})|$ and $h_{K}^{-}=h_{K}/h_{K}^{+}$ (which is an integer). Let $\operatorname{disc}(K)$ denote the discriminant of a number field $K$.
@@
We use that, if $L/K$ is an extension of abelian fields, then $\operatorname{odd}(h_{K}^{-})\mid\operatorname{odd}(h_{L}^{-})$ [Ho89, Lemma 1]. For each abelian field $K$, let $K^{\prime}$ denote the maximal subfield of $K$ with degree a power of 2. By the fundamental theorem of Galois theory and the fact that a finite abelian group $A$ has a subgroup of order $d$ for all $d\mid|A|$, we get that $|K^{\prime}/\mathbb{Q}|$ is the highest power of 2 dividing $|K/\mathbb{Q}|$. For an integer $n\geq 1$, let $\mathcal{A}_{n}=\{m\mid\operatorname{odd}(h_{m}^{-})\leq n\}$ and $\mathcal{B}_{n}=\{|\mathbb{Q}(\zeta_{m})^{\prime}/\mathbb{Q}|\mid m\in\mathcal{A}_{n}\}$.
We begin by finding a bound for $\sup(\mathcal{B}_{n})$. Let $K=\mathbb{Q}(\zeta_{m})^{\prime}$ for some $m\in\mathcal{A}_{n}$. Then $\mathbb{Q}(\zeta_{m})/K$ is an extension of abelian fields and so $\operatorname{odd}(h_{K}^{-})\mid h_{m}^{-}$ and so $\operatorname{odd}(h_{K}^{-})\leq n$. Since $|K/\mathbb{Q}|$ is a power of 2, $h_{K}$ must be odd [Wa97] and so $h_{K}^{-}=\operatorname{odd}(h_{K}^{-})\leq n$. Furthermore, $K$ is imaginary unless $K=\mathbb{Q}$ [Ho89, p. 468].
@@
Let $K=\mathbb{Q}(\zeta_{m})^{\prime}$. Then $|\mathbb{Q}(\zeta_{m})/\mathbb{Q}|=2^{r}t$ for some $r\geq 0$ and $t$ odd, where $2^{r}=|K/\mathbb{Q}|\leq C\log n$ by the bound on $\mathcal{B}_{n}$. Since $|\mathbb{Q}(\zeta_{m})/L|\mid|\mathbb{Q}(\zeta_{m})/\mathbb{Q}|$, we have that $2^{n(2)}\leq 2^{r}\leq C\log n$. Hence $h_{m}^{-}/\operatorname{odd}(h_{m}^{-})\leq h_{m}/\operatorname{odd}(h_{m})=2^{e_{N}^{(2)}}\leq(C\log n)^{C^{\prime}}$. Since $m\in\mathcal{A}_{n}$, we have $\operatorname{odd}(h_{m}^{-})\leq n$. This gives that $h_{m}^{-}\leq a(\log n)^{b}n$ for some constants $a,b>0$ and so, for any $\varepsilon>0$, we have that $h_{m}^{-}\leq an^{1+\varepsilon}$. Combining this with Proposition 10.7 gives that $\log(an^{1+\varepsilon})\geq C_{0}\frac{m\log m}{\log\log m}$ for some $C_{0}>0$, which implies that $\log n\geq C\frac{m\log m}{\log\log m}$ for some $C>0$.
@@
We now consider the parity of $h_{m}$ and $h_{m}^{-}$. Whilst we do not explicitly make use of it, we record the following basic observation which dates back to Kummer (see [Ha52, Satz 45], [Yo98, Remark 1]).
@@
Prior to the results of Ichimura–Nakajima, it was shown by Washington that $h_{p^{n}}/h_{p}$ is odd for $p=3,5$ [Wa75]. Both results of Washington and Ichimura–Nakajima all depend on Iwasawa theory. As far as we are aware, this is an essential ingredient in all known proofs that there exists an odd prime $p$ such that $h_{p^{n}}$ is odd for all $n\geq 1$.
@@ 10.4. Kernel groups of $\mathbb{Z}C_{m}$
The aim of this section is to determine the involution on $D(\mathbb{Z}C_{m})$ which is induced by the involution on $C(\mathbb{Z}C_{m})$ (see Proposition 8.12).
We begin with the following classical result due to Rim [Ri59, Theorem 6.24] (see also [CR87, Theorem 50.2]). Recall that, if $p$ is a prime and $\Gamma_{p}$ is the maximal order in $\mathbb{Q}C_{p}$ containing $\mathbb{Z}C_{p}$, then $\Gamma_{p}\cong\mathbb{Z}\times\mathbb{Z}[\zeta_{p}]$ and so $C(\Gamma_{p})\cong C(\mathbb{Z}[\zeta_{p}])$ since $\mathbb{Z}$ is a PID.
@@
We now determine $D(\mathbb{Z}C_{m})$, as well as its involution, in the case where $m$ is square-free. This is used in the proof of Theorem 11.1. As usual, we view an abelian group with involution as a $\mathbb{Z}C_{2}$-module.
@@
We view this as a $\mathbb{Z}C_{2}$-module with the involution induced by the conjugation map
$\overline{\cdot}\colon(\mathbb{Z}[\zeta_{m}]/\pi_{m})^{\times}\to(\mathbb{Z}[\zeta_{m}]/\pi_{m})^{\times},\quad\zeta_{m}\mapsto\zeta_{m}^{-1}.$
If $p$ is prime, then $\mathbb{Z}[\zeta_{p}]/\pi_{p}\cong\mathbb{F}_{p}$. We can then see that $\Psi_{p}\colon\mathbb{Z}[\zeta_{p}]^{\times}\to\mathbb{F}_{p}^{\times}$ is surjective by considering the cyclotomic units $1+\zeta_{p}+\cdots+\zeta_{p}^{i-1}\in\mathbb{Z}[\zeta_{p}]^{\times}$ for $(i,p)=1$. In particular, $V_{p}=1$. We now show how $D(\mathbb{Z}C_{m})$ is related to $V_{d}$ for $d\mid m$.
@@
Firstly, $\Gamma=\bigoplus_{d\mid m}\mathbb{Z}[\zeta_{d}]$ is the unique maximal order in $\mathbb{Q}C_{m}$ which contains $\mathbb{Z}C_{m}$. This implies that there is a pullback square:
$\mathbb{Z}C_{m}$$\Gamma$$\mathbb{Z}C_{m}/m\Gamma$$\Gamma/m\Gamma$$\scriptstyle{i_{2}}$$\scriptstyle{i_{1}}$$\scriptstyle{j_{2}}$$\scriptstyle{j_{1}}$
where the inclusion $m\Gamma\subseteq\mathbb{Z}C_{m}$ follows from [Re75, Theorem 41.1].
@@
is the $\mathbb{Z}C_{m}$-module with action $\lambda\cdot(x,y)=(i_{1}(\lambda)x,i_{2}(\lambda)y)$ for $\lambda\in\mathbb{Z}C_{m}$. We now claim that the conjugation map on $(\Gamma/m\Gamma)^{\times}$ induces the involution on $D(\mathbb{Z}C_{m})$. Firstly, by Proposition 8.13, we have that the involution on $D(\mathbb{Z}C_{m})$ is induced by the natural involution $x\mapsto x^{-1}$ on the idèle group $J(\mathbb{Q}C_{m})\subseteq\prod_{p}\mathbb{Q}_{p}C_{m}$. For all primes $p$, we have $\mathbb{Z}_{p}C_{m}\subseteq\Gamma_{p}\oplus(\mathbb{Z}C_{m}/m\Gamma)_{p}\subseteq\mathbb{Q}_{p}C_{m}$. If $M(u)=\mathbb{Z}C_{m}\alpha$ for $\alpha=(\alpha_{p})\in J(\mathbb{Q}C_{m})$, then [CR87, Exercise 53.1] implies that
@@
where $u_{p}\in\Gamma_{p}$ is any element such that $j_{2}(u_{p})=[u]\in(\Gamma/m\Gamma)_{p}$. By the same argument as in the proof of Proposition 10.1, the involution on $J(\mathbb{Q}C_{m})$ induces an involution on $J(\mathbb{Q}(\zeta_{d}))$ which coincides with the involution induced by conjugation. In particular, the involution maps $\alpha_{p}\mapsto(1,1)$ or $(\overline{u}_{p},1)$ where $\overline{\,\cdot\,}\colon\Gamma_{p}\to\Gamma_{p}$ is induced by conjugation on $\Gamma$. In particular, this coincides with the involution induced by conjugation on $(\Gamma/m\Gamma)^{\times}$.
By [CR87, Lemma 50.7], $j_{1}$ can be replaced by the map $\alpha\colon(\mathbb{Z}C_{m}/m\mathbb{Z}C_{m})^{\times}\to(\Gamma/m\Gamma)^{\times}$. By [CR87, Lemma 50.8], $\operatorname{coker}(\alpha)\cong\bigoplus_{d\mid m}(\mathbb{Z}[\zeta_{d}]/\pi_{d})^{\times}$ and it follows from the proof that the conjugation map on $(\Gamma/m\Gamma)^{\times}$ induces conjugation on $(\mathbb{Z}[\zeta_{d}]/\pi_{d})^{\times}$ for each $d\mid m$. If $\gamma\colon\Gamma^{\times}\to\operatorname{coker}(\alpha)$ is the map induced by $j_{2}$, then we obtain an exact sequence
$\bigoplus_{d\mid m}\mathbb{Z}[\zeta_{d}]^{\times}\xrightarrow[]{\gamma}\bigoplus_{d\mid m}(\mathbb{Z}[\zeta_{d}]/\pi_{d})^{\times}\xrightarrow[]{\overline{\partial}}D(\mathbb{Z}C_{m})\to 0.$
@@
Since the involution on $D(\mathbb{Z}C_{m})$ is induced by conjugation on $(\Gamma/m\Gamma)^{\times}$, it follows that it restricts to $A_{i}$ where it acts via conjugation on the $\mathbb{Z}[\zeta_{d}]$. Hence, with respect to the involution, $A_{i}\leq D(\mathbb{Z}C_{m})$ is a $\mathbb{Z}C_{2}$-module and the chain $A_{0}\leq\cdots\leq A_{n}$ is a chain of $\mathbb{Z}C_{2}$-modules. Under the abelian group isomorphism
@@
Let $m\geq 2$ be a square-free integer. We now give a method for analysing the involution on $V_{m}=\operatorname{coker}(\Psi_{m})$. For a field $\mathbb{F}$ and $\alpha_{1},\dots,\alpha_{n}\in\mathbb{C}$, we write $\mathbb{F}[\alpha_{1},\dots,\alpha_{n}]$ to denote $\mathbb{F}\otimes_{\mathbb{Z}}\mathbb{Z}[\alpha_{1},\dots,\alpha_{n}]$. Note that this may not be a field and so need not coincide with $\mathbb{F}(\alpha_{1},\dots,\alpha_{n})$.
Firstly, as described above, we have that
@@
We view this as a $\mathbb{Z}C_{2}$-module with the involution induced by the conjugation map
$\overline{\cdot}\colon\mathbb{F}_{p_{i}}[\zeta_{m/p_{i}}]^{\times}\to\mathbb{F}_{p_{i}}[\zeta_{m/p_{i}}]^{\times},\quad\zeta_{m/p_{i}}\mapsto\zeta_{m/p_{i}}^{-1}.$
@@
and $f(\overline{\alpha})=f(\alpha)^{-1}$. Since $f$ is surjective and induces the involution on $\widetilde{V}_{m}$, this implies that $\overline{x}=-x$ for all $x\in\widetilde{V}_{m}$, where we now write the inverse as x-x rather than $x^{-1}$ since $\widetilde{V}_{m}$ is an abelian group. ∎
We now deduce the following, which is the main result of this section. This is analogous to Proposition 10.5 which applied in the case of ideal class groups. Recall that, for an integer $m$, we let $\operatorname{odd}(m)$ denote the unique odd integer $r$ such that $m=2^{k}r$ for some $k$.
@@
In particular, $\prod_{d\mid m}\operatorname{odd}(|\widetilde{V}_{d}|)$ divides $|\{x-\overline{x}\mid x\in D(\mathbb{Z}C_{m})\}|$.
@@
and $|2\cdot\widetilde{V}_{d_{n-i+1}}|$ divides $|\{x-\overline{x}\mid x\in A_{i}\}/\{x-\overline{x}\mid x\in A_{i-1}\}|$ for all $1\leq i\leq n$. It now suffices to observe that $\operatorname{odd}(|\widetilde{V}_{d_{n-i+1}}|)$ divides $|2\cdot\widetilde{V}_{d_{n-i+1}}|$ by the same argument as in the proof of Proposition 10.5. ∎
@@
To obtain a complete analogue of Proposition 10.5, it would be desirable to also obtain bounds on $|\{x\in D(\mathbb{Z}C_{m})\mid x=-\overline{x}\}|$. However, unlike Proposition 10.5, the bounds we obtain in Proposition 10.18 are obtained by subquotients rather than just subgroups. We therefore cannot apply Lemma 9.3 (i) since the final map in the sequence need not be surjective in general. As we shall see in Section 11, it is possible to circumvent the need for such bounds in the proof of Theorem 11.1.
We now conclude this section with a result which holds in the case that $m$ is not square-free. Firstly an analogue of Lemma 10.15 holds in the case that $m$ is a prime power, by Kervaire–Murthy [KM77, Theorem 1.2]. For brevity, we do not state this result here. We instead make do with the following consequence of their result in the case that $p=2$ which is used in the proof of Theorem 11.2. Note that $V_{2^{n+1}}$ is directly analogous to the $\mathbb{Z}C_{2}$-modules $V_{m}$ defined in Definition 10.14 when $m$ is square-free.
@@
$0\to V_{2^{n+1}}\to\widetilde{K}_{0}(\mathbb{Z}C_{2^{n+1}})\xrightarrow[]{\alpha}\widetilde{K}_{0}(\mathbb{Z}C_{2^{n}})\oplus\widetilde{K}_{0}(\mathbb{Z}[\zeta_{2^{n+1}}])\to 0$